OpenStudy (anonymous):

the lenght of a certain rectangle is six more than three time its widht. if the width of the rectangle is 4 units,wat is its lenght?

OpenStudy (anonymous):

L= 3W+6 L=3(4)+6 L=12+6 L=18

OpenStudy (anonymous):

thanks